OpenStudy (ana.ie.lee):

how do you graph y=-x+4, please show both points

OpenStudy (sunnnystrong):



OpenStudy (ana.ie.lee):

what are the coordinates for the second one

OpenStudy (ana.ie.lee):

i know the first is (0,4) but whats the second coordinate

jhonyy9 (jhonyy9):

just you need give to x two values and calculi the y so than you need graph these two point with coordinates what will get hope helped

OpenStudy (sunnnystrong):

well.. if y=mx+b and your looking at the linear equation of y=-x+4 Than the y-int is (0,4) and you can find another point on the graph using the slope m=change in y over change in x so the next point would be 1,3 and so forth...

OpenStudy (sunnnystrong):

you can also just plug in values for x(input) and see what values of y you get(output) than graph those

OpenStudy (ana.ie.lee):

can you also help me graph y=2x-5

OpenStudy (sunnnystrong):

@ana.ie.lee sure! the y intercept is when x=0 so... y=2(0)-5 the y-intercept is -5. Finding a second point on the graph... you can use x=1 when x=1 y=2(1)-5 y= -3 Your two points : (0,-5) and (1,-3) All you need to do is graph

OpenStudy (sunnnystrong):